# C2SaferRust: Transforming C Projects into Safer Rust with NeuroSymbolic Techniques

### v1

@@ 7 Threats to Validity
Our metrics are intended to model the safety and naturalness of the code; however it is possible that an improvement in a metric doesn’t correlate with improved safety or naturalness of the code. For example, the number of lines of code within unsafe blocks could decrease simply by eliminating line breaks and relying on semicolons to delimit lines. We therefore assume that the tools we consider do not attempt to game the metrics by performing such manipulations.
@@ 8 Conclusion
In this work, we have presented C2SaferRust, a system to automatically translate full C repositories to safer Rust. Based on the novel insight that the C2Rust transpiler can be used as an intermediary to bridge between C and Rust for the purpose of LLM-aided translation, we first convert the entire repository into unsafe Rust. We propose an algorithm to divide the code repository into smaller “translation units”, and extract each translation unit along with its context information. We use an LLM to translate each unit to safer Rust, swap in this translation, and propagate the changes to the rest of the code base.
We contribute a dataset of 7 C programs from GNU CoreUtils with their corresponding C2Rust transpiled versions. Each program comes with end-to-end test cases, allowing one to empirically verify the correctness of translation. On this dataset, we show that we reduce raw pointer declarations and dereferences by up to 38%, and reduce the amount of unsafe code by up to 28%. We also compare our tool with two baselines for making Rust code safer, and observe consistent gains. Future work could look at translating code outside functions, like structure definitions, as well as remapping C FFI calls to safe Rust equivalents.
@@ 14 Block => visit_block(child),
31 while upper $<$ block.statements.len():
@@
43 if lower $<$ block.statements.len():

### v2

@@ 7 Threats to Validity
Although C2Rust makes efforts to ensure functional equivalence between the C and unsafe Rust code, this is not a formal guarantee. C2SaferRust uses C2Rust as the first step, and so the validity of the entire approach depends on the correctness of C2Rust’s transpilation. However, C2Rust merely provides an interface between the original code and translated code that enables the program to be tested end-to-end. It could be replaced with any other such framework, like Rust’s Foreign Function Interface (FFI).
Our metrics are intended to model the safety and naturalness of the code; however it is possible that an improvement in a metric doesn’t correlate with improved safety or naturalness of the code. For example, the number of lines of code within unsafe blocks could decrease simply by eliminating line breaks and using semicolons to delimit lines. We therefore assume that the tools we consider do not attempt to game the metrics by performing such manipulations.
@@ 8 Conclusion
In this work, we have presented C2SaferRust, a system to automatically translate full C repositories to safer Rust. Based on the novel insight that the C2Rust transpiler can be used as an intermediary to bridge between C and Rust, we first convert the entire repository into unsafe Rust. We divide the code repository into smaller “translation units”, and extract each translation unit along with its context information. We use an LLM to translate each unit to safer Rust, swap in this translation, and propagate the changes to the rest of the code base. We contribute a dataset of 7 C programs from GNU CoreUtils with their corresponding C2Rust transpiled versions. Each program comes with end-to-end test cases, allowing one to empirically verify the correctness of translation. On this dataset, we show that we reduce raw pointer declarations and dereferences by up to 38%, and reduce unsafe code by up to 28%. We also compare our tool with two baselines for making Rust code safer, and observe consistent gains. Future work could look at translating code outside functions, like structure definitions, as well as remapping C FFI calls to safe Rust equivalents.
@@ 14 Block => visit_block(child),
31 while upper << block.statements.len():
@@
43 if lower << block.statements.len():
@@
Appendix B More Coreutils Results
In this section, we present some more results on the Coreutils datasets, adding to the results in Table˜II. Table˜V contains the results of a variant of our approach that uses direct C to Rust translation. Table˜VI shows the number of LLM calls and approximate usage cost for all our benchmarks.
TABLE V: Translating CoreUtils to safer Rust using a variant of C2SaferRust, with direct C to Rust translation.
    Raw Pointer     Raw Pointer
    Declarations ($\nabla$)     Deferences ($\nabla$)
Before
After
$\Delta\%$
Before
After
$\Delta\%$
     split     252    232    8    656    628    4
   pwd     164    138    16    295    265    10
   cat     192    163    15    317    289    9
   truncate     156    142    9    326    278    15
   uniq     227    179    21    343    310    10
   tail     389    338    13    1092    1027    6
   head     192    172    10    442    420    5
(a) Number of raw pointer declarations and dereferences.
    Unsafe     Unsafe     Unsafe Call
    Lines ($\nabla$)     Type Casts ($\nabla$)     Expressions ($\nabla$)
Before
After
$\Delta\%$
Before
After
$\Delta\%$
Before
After
$\Delta\%$
     split     11324    10357    9    2353    2326    1    11324    10357    9
   pwd     4201    2921    30    875    851    3    4201    2921    30
   cat     5625    4663    17    1038    950    8    5625    4663    17
   truncate     5544    4581    17    1040    1050    -1    5544    4581    17
   uniq     6066    4986    18    1150    1129    2    6066    4986    18
   tail     11663    9780    16    2580    2485    4    11663    9780    16
   head     6245    5552    11    1378    1383    -0    6245    5552    11
(b) Number of unsafe expressions and declarations.
Program # Fn
# Trans.
Units
LLM
Calls
Approx
Cost (\$)
Coreutils
split 207 292 1140 0.68
pwd 127 152 527 0.31
cat 166 205 636 0.38
truncate 124 170 599 0.36
uniq 167 213 776 0.46
tail 266 350 1224 0.73
head 153 193 666 0.40
Laertes
tulipindicators 234 242 1017 0.61
xzoom 11 26 101 0.06
genann 32 35 124 0.07
optipng 576 941 3518 2.10
urlparser 22 25 101 0.06
lil 160 184 800 0.48
qsort 3 3 12 0.01
snudown 92 108 488 0.29
grabc 7 7 22 0.01
bzip2 128 251 885 0.53
TABLE VI: LLM usage cost estimates for translating each benchmark using C2SaferRust with gpt4o-mini.
Appendix C Prompts
Here is a function:
‘‘‘rust
unsafe extern "C" fn gettext_quote(
mut msgid: *const libc::c_char,
mut s: quoting_style,
) -> *const libc::c_char {
let mut translation: *const libc::c_char = gettext(msgid);
let mut locale_code: *const libc::c_char = 0 as *const libc::c_char;
if translation != msgid {
return translation;
locale_code = locale_charset();
Here are its call sites
Call site 1:
‘‘‘rust
right_quote = gettext_quote(
b"’\0" as *const u8 as *const libc::c_char,
quoting_style,
Call site 2:
‘‘‘rust
left_quote = gettext_quote(
b"‘\0" as *const u8 as *const libc::c_char,
quoting_style,
The file contains the following imports:
‘‘‘rust
use ::libc;
Convert the function to idiomatic Rust, meaning Rust code that does not make use of features like unsafe, raw pointers, and the C API whenever possible. Do not change the function name.
Follow the following format for your output: Place the function translation between the tags <FUNC> and </FUNC>. If the function signature changed in translation, its callsites will need to be modified as well. Place each callsite translation (in the same order it appears above) between <CALL> and </CALL>. Note that even if the callsite is only a single statement, the translation can be mutiple statements. For example, you may need to declare new variables, or convert between types, either before or after the call. The translation should be such that the surrounding code is not affected by the changes.
Any functions or variables without definitions are defined elsewhere in the code. Do not attempt to redefine them or import them.
If you are using any new functions and you need new imports for those, place them between the tags <IMPORTS> and </IMPORTS>. This should be only *new* imports. Do not include existing imports.
DO NOT include markdown characters like "‘‘‘" or "‘‘‘rust" in your translation.
TABLE VII: Translation prompt for a Coreutils function.
Appendix D Test Statistics
Table˜VIII shows the statistics of the Coreutils test suite that we used to evaluate the correctness of our translations. Each test is a Bash or Perl script that is run against the compiled executable.
Program # Fn
# Cov.
Fn
# Test
Scripts
Coreutils
split 207 73 12
pwd 127 16 2
cat 166 37 4
truncate 124 33 8
uniq 167 34 3
tail 266 76 30
head 153 45 4
TABLE VIII: The statistics of the test suite for the 7 programs in our coreutils benchmark dataset. “# Cov. Fn” is the number of functions covered by tests, and “# Test Scripts” is the number of end-to-end test scripts.
Appendix E More Qualitative Comparisons
In this section, we show some more examples of C2SaferRust output that will enable us to qualitatively analyze its performance.
Figure˜7 shows an example of the output C2Rust vs Laertes vs C2SaferRust. This is taken from qsort, one of the Laertes benchmarks. We can see that although both Laertes and C2SaferRust eliminate the raw pointer *mut, the output of C2SaferRust is much more natural and easy to follow.
pub unsafe extern "C" fn swap(
mut a: *mut std::os::raw::c_int,
mut b: *mut std::os::raw::c_int) {
let mut t: std::os::raw::c_int = *a;
*a = *b;
*b = t;
(a) C2Rust output
pub unsafe extern "C" fn swap<’a1, a2>(
mut a: Option<&’a1 mut std::os::raw::c_int>,
mut b: Option<&’a2 mut std::os::raw::c_int>) {
let mut t: i32 = *(borrow_mut(&mut a)).unwrap();
*(borrow_mut(&mut a)).unwrap() =
*(borrow_mut(&mut b)).unwrap();
*(borrow_mut(&mut b)).unwrap() = t;
(b) Laertes output.
pub fn swap(a: &mut i32, b: &mut i32) {
let t = *a;
*a = *b;
*b = t;
(c) C2SaferRust output.
Figure 7: A qualitative comparison of Laertes and C2SaferRust.
Figure˜8 shows the function ireallocarray, which is a common helper function used by all the Coreutils programs. It takes a pointer pointing to potentially uninitialized memory, and allocates memory for it such that the total allocation is equal to a certain specified number of bytes. This is implemented in C with a void * pointer. The C2Rust code uses *mut libc::c_void, which is not an idiomatic Rust type. On the other hand, C2SaferRust correctly handles this by using the type Option<Vec<u8>> for p. This also allows for the possibility that the memory might be uninitalized, which would correspond to the None variant of the Option enum.
/* ireallocarray (ptr, num, size) is like reallocarray (ptr, num, size).
It returns a non-NULL pointer to num * size bytes of memory.
Upon failure, it returns NULL with errno set. */
IALLOC_INLINE void *
ireallocarray (void *p, idx_t n, idx_t s)
if (n <= SIZE_MAX && s <= SIZE_MAX)
size_t nx = n;
size_t sx = s;
if (n == 0 $|$ s == 0)
nx = sx = 1;
p = reallocarray (p, nx, sx);
return p;
else
return _gl_alloc_nomem ();
(a) C code
pub unsafe extern "C" fn ireallocarray(
mut p: *mut libc::c_void,
mut n: idx_t,
mut s: idx_t,
) -> *mut libc::c_void {
if n as libc::c_ulong <= 18446744073709551615
as libc::c_ulong
&& s as libc::c_ulong <= 18446744073709551615
as libc::c_ulong
let mut nx: size_t = n as size_t;
let mut sx: size_t = s as size_t;
if n == 0 as libc::c_int as libc::c_long
$|$ s == 0 as libc::c_int as libc::c_long
sx = 1 as libc::c_int as size_t;
nx = sx;
p = reallocarray(p, nx, sx);
return p;
} else {
return _gl_alloc_nomem()
(b) C2Rust output.
pub fn ireallocarray(
p: Option<Vec<u8>>,
n: idx_t,
s: idx_t,
) -> Option<Vec<u8>> {
if n as usize <= usize::MAX && s as usize <= usize::MAX {
let mut nx: usize = n as usize;
let mut sx: usize = s as usize;
if n == 0 $|$ s == 0 {
sx = 1;
nx = sx;
let mut vec = p.unwrap_or_else($|$
Vec::with_capacity(nx * sx)
vec.resize(nx * sx, 0);
return Some(vec);
} else {
return None;
(c) C2SaferRust output.
Figure 8: A comparison between C, C2Rust, and C2SaferRust, for an example involving memory allocation. C2SaferRust appropriately handles the allocation using Option<Vec<u8>> in Rust.